# Entropy and forecasting complexity of hidden Markov models, matrix product states, and observable operator models

### v1

@@ Abstract
In a series of three papers, Jurgens and Crutchfield recently proposed a supposedly novel method to compute entropies of hidden Markov models (HMMs), discussed in detail its relationship to iterated function systems, and applied it to compute their “ambiguity rates”, a concept supposedly introduced by Claude Shannon. We point out that the basic formalism is not new (it is the well known “forward algorithm” for HMMs), and that all three papers have also serious other faults.
@@
In Jurgens1 , a supposedly new algorithm based on “mixed states” was proposed. But this algorithm is exactly the forward algorithm, and the “mixed states” become precisely the causal states introduced in Grass86 , when their set is minimized. This minimization is, in general, anyhow necessary for an efficient application. Indeed, the forward algorithm was independently (re-)discovered for a special class of HMMs in Zambella , where it was used to present precise entropy estimates for a number of non-trivial models, and where it was also used to compute forecasting complexities (FCs Grass86 ; later called ‘stochastic complexities’ in Young ). The fact that the algorithm proposed in Zambella is much more general was stressed in arXiv , where also a number of other points were clarified.
More precisely, we consider a finite stationary ergodic irreducible first order Markov process $M$. Here, “finite” means that it has a finite number $N$ of states, in the following called $s_{t}$ with values numbered simply by ``$i",\;i=1\ldots N$. The restriction to first order is trivial, as any higher order Markov chain can be re-written as a first order process by re-labelling the states. Irreducibility is essential since it implies (by the Frobenius-Perron-theorem) that there is a unique stationary state. The probability to be in state $i$ at (discrete) time $t$ is denoted as $P_{i,t}$. The stationary state is characterized by the probabilities $P_{i,\infty}$. The transition matrix is $T_{ji}={\rm prob}\{s_{t+1}=j|s_{t}=i\}$. The time evolution of the probabilities is
@@
The optimally inferred states of the hidden dynamics based on the observation process were called ‘mixed states’ in Jurgens1 ; Jurgens2 ; Jurgens3 , and will be denoted here as $\hat{\bf P}_{t}(x_{1}x_{2}\ldots x_{t})$. Based on them, the optimal forecast of the observed process is $\hat{p}_{t+1}(x|x_{1}x_{2}\ldots x_{t})=\hat{p}_{t+1}(x_{1}x_{2}\ldots x_{t},x)/\hat{p}_{t}(x_{1}x_{2}\ldots x_{t})$ with
@@
Equations (4) and (6) are essentially the same as Eq.(3.4) in Zambella , and are today well known as ‘forward algorithm’ Zucchini . They hold also for ‘observable operator models’ (OOMs) Jaeger ; Spanczer (in which the positivity requirements made for HMMs are dropped), and even for ‘matrix product states’ in statistical physics Derrida ; Klumper ) with discrete time and discrete values.
@@
As shown in Zambella (see also arXiv ), the set of ‘mixed states’ becomes precisely the set of causal states introduced in Grass86 when minimized. Minimization means that two vectors $\hat{\bf P}_{t}(x_{1}\ldots x_{t})$ and $\hat{\bf P}_{t}(x^{\prime}_{1}\ldots x^{\prime}_{t})$ are identified if they lead, via Eq.(4), to the same $\hat{p}_{t+k}(x)\;\forall k\geq 1$. No such minimization was done in Jurgens1 , which is why no FCs were estimated there and no precise estimates of entropy. It was also shown in Zambella that causal states do in general not correspond to elements of a partitioning of phase space (in contrast to what was claimed in Young ), at least when transient parts of the “$\epsilon$-machine” are not cut off. Cutting them off (as proposed in most recent papers by Crutchfield et al.) is not a viable alternative, as in some cases (called ‘totally recurrent graphs’ in arXiv ) none of the nodes of the “$\epsilon$-machine” correspond to elements of a phase space partitioning Zambella ; arXiv . An example of the latter is e.g. ‘rule 22’ Zambella ; arXiv .
@@
A main point in Jurgens1 ; Jurgens2 was that no such minimization can in general be expected (and FCs are infinite), if the HMM depends on irrational real valued parameters, such that all different observed sequences lead to different optimal forecasts. It was already pointed out in Grass-Helv ; Grass-Nat that even rather simple systems can lead to infinite FC. On the other hand, symmetries of the ${\bf T}^{(x)}$ can lead to non-trivial minimizations, even if the ${\bf T}^{(x)}$ depend on irrational parameters (an instance where some optimal forecasts do not depend on the history at all, but FC is infinite nevertheless, will be be discussed below). No entropy estimates with precision close to these in Zambella were made in Jurgens1 , and no estimates of FCs at all. The main reason is that no minimizations of the set of ‘mixed states’ were performed.
@@
It is claimed in Jurgens1 ; Jurgens2 ; Jurgens3 that the set of mixed states is in general not countable, if irrational parameters are involved, but this is wrong. It is always countable by construction. What is correct, however, is that the closure of the set of mixed states is in such cases uncountable, just as the triadic Cantor set is countable, while its closure is not (a non-countable set could be obtained by using any arbitrary starting vector ${\bf P}_{0}$ instead of ${\bf P}_{0}={\bf P}_{\infty}$, but the resulting forecasts would not be optimal). This non-countability of the closure is, however, not restricted to HMMs with real-valued parameters (as claimed in Jurgens1 ; Jurgens2 ; Jurgens3 ), but holds also for many of the HMMs studied in Zambella . The difference between these two groups of models is not in the cardinality of the sets of inferences, but in the Hausdorff dimension of the invariant measure (the Blackwell measure Blackwell ) supported by it. It was zero in all models studied in Zambella , although the box-counting dimensions were non-zero in many cases.
@@
The failure of distinguishing sets from their closures is also responsible for most of Jurgens2 being obsolete. As remarked already in Zambella , the set of mixed (=causal) states is very similar to an iterated function system (IFS), with a difference largely overlooked in Jurgens1 ; Jurgens2 ; Jurgens3 : While the maps ${\bf T}^{(x)}$ are chosen randomly (maybe with non-uniform probabilities) in an IFS, their choice is in the present case constrained by the (in general non-trivial) grammar of the sequence $\{x_{t}\}$. Thus theorems about IFSs cannot be taken over blindly  [^1] Non-trivial grammatical restrictions do not seem to be responsible for the ‘resetting strings’ found in some of the models studied in Zambella , which are finite strings $R=x_{1}\ldots x_{n}$ of observed symbols for which ${\bf T}^{(R)}\equiv{\bf T}^{(x_{n})}\ldots{\bf T}^{(x_{1})}$ maps any ${\bf P}$ onto ${\bf P}_{\infty}$. They correspond to catastrophic forgettings of the internal state and are then responsible for the finite Hausdorff dimension of the Blackwell measure; in a systematic search, we now found that this phenomenon occurs in 108 of the 256 models studied in Zambella . For the model called ‘rule 22’ in Zambella . the shortest resetting string is e.g. ‘01000010000’ (the string given in Zambella is an extension of this).. Nevertheless, the open set condition (OSC) Moran ; Peres-Rams should hold in many cases. But its application is much less trivial than assumed in Jurgens2 , where the (non-)overlapping of the sets was confused with the (non-)overlapping of their closures. That such a confusion can be fatal is maybe best illustrated by the ‘fat baker’s transformation’ Alexander .
@@
If minimization of the set of mixed states had been used, the study of the Hausdorff dimension of the set of causal states in Jurgens2 (called ‘statistical complexity dimension’ there), would have been very useful. Both were obviously motivated by the calculations of Hausdorff dimensions and by similar studies of the convergence of entropy and forecasting complexity in Zambella , although that paper was not cited. But although this minimization was alluded to repeatedly in Jurgens1 ; Jurgens2 ; Jurgens3 , it was never actually done.
@@
Indeed, this can be defined for general stationary ergodic processes, not only for HMM. All what is needed is that one can define a sequence of approximate HMMs or OOMs, to which one can apply Eqs. (4,6). Minimizing $h_{C}-h$ can be used for model (or parameter) inference Zucchini , but minimizing both $h_{C}-h$ and $C$ at fixed $\alpha$ could be useful for finding models / parameters / approximations that are both accurate and relatively easy to forecast.
@@
Finally, Ref. Jurgens3 is devoted to the calculation of the ‘ambiguity rate’ $h_{a}$ of a process, a notion supposedly introduced by Shannon in Shannon as a quantity opposed to ‘equivocation’. In fact, ‘ambiguity’ is mentioned exactly once in Shannon , but as a synonym for equivocation. The concept as defined in Jurgens3 is the rate at which information about the past can be discarded, and yet prediction of the future is optimal. If $h_{a}<h$, this would mean that less information is dismissed than new information is obtained, i.e. that FC (the information stored in an optimal forecasting scheme) diverges with time. In the opposite case, i.e. when $h_{a}=h$, it was claimed that “the associated $\epsilon$-machine’s internal causal-state process is stationary.” Jurgens3 . This is wrong, as demonstrated by the counter-examples given in Zambella . There it was shown that FC is finite in all considered cases (i.e., $h_{a}=h$), but yet the forecasting process is not stationary in cases like ‘rule 22’, because of the slow convergence of FC.
@@
When each observed sequence leads to a different forecast, it is clear that no information can be dismissed in an optimal forecast, and thus $h_{a}=0$. In some cases, the situation is a bit more subtle. Consider, e.g., the example in Sec. 7B of Jurgens3 for control parameter values $x=0.25,\alpha=0.5$, which was treated in detail in appendix B of Jurgens3 . In this model, there are three emitted symbols ,$\triangle$,$\Box$,$\circ$ , □ , ∘. One finds that always $\hat{p}_{t}(\Box)=1/4$, while $\hat{p}_{t}(\triangle)$ and $\hat{p}_{t}(\circ)$ depend non-trivially on the observations at times $11,\ldots t-$ 1. Indeed, no two different histories lead to the same $\hat{p}_{t}(\triangle)$ and $\hat{p}_{t}(\circ)$ (at least for $t<12$; for larger $t$ one would need extended precision arithmetic to avoid numerical problems), so that no information can be dismissed if all forecasts are to be optimal, and thus $h_{a}=0$ — while $h_{a}=0.4499$ bits is quoted in Jurgens3 . Obviously, this wrong value was obtained because $h_{a}$ was calculated in Jurgens3 using the open set condition, and overlaps between sets were confused with overlaps between their closures. Since this error was made in all calculations of $h_{a}$, all numerical results of Jurgens3 are either trivial or obsolete.
In summary, we have shown that most results of Jurgens1 ; Jurgens2 ; Jurgens3 are either wrong or were well known for a long time. In particular, many of them were first obtained in Zambella (which was not cited), a fact pointed out already several times to Dr. Crutchfield (see also arXiv ; arXiv2 ).

### v2

# On three papers by Jurgens & Crutchfield, and on the basic structure of “computational mechanics”

@@ Abstract
In a recent paper, Jurgens and Crutchfield [Phys. Rev. E 104, 064107 (2021), called “paper III” in the following] computed what they called the “ambiguity rate” of hidden Markov processes, a concept supposedly introduced by Claude Shannon. This calculation was based on a “mixed state” formalism introduced by them in J. Stat. Phys. 183, 32 (2021) (“paper I”), and developed further in Chaos, 31, 083114 (2021) (“paper II”). We point out that (i) ambiguity rates were not introduced by Shannon; (ii) their computations in paper III are wrong, because of an error made already in papers I and II; (iii) due to this error (a confusion between open sets and their closures), also many of the “statistical complexity dimensions” computed in II are wrong; (iv) the “mixed state” formalism of I is just the well known ‘forward algorithm’ for hidden Markov models; (v) the ‘causal states’ in $\epsilon$-machines’ correspond in general to finite (as opposed to infinite, as often claimed) histories; and (vi) $\epsilon$-machines’ are always countable, in contrast to frequent claims in the literature. In addition, we propose an alternative complexity measure for models where the forecasting complexity is infinite, and we point out that our results apply also beyond hidden Markov models.
@@
In Jurgens1 (paper I), a supposedly new algorithm based on “mixed states” was proposed. But this algorithm is exactly the forward algorithm, and the “mixed states” become precisely the causal states introduced in Grass86 , when their set is minimized. This minimization is, in general, anyhow necessary for an efficient application. Indeed, the forward algorithm was independently (re-)discovered for a special class of HMMs in Zambella , where it was used to present precise entropy estimates for a number of non-trivial models, and where it was also used to compute forecasting complexities (FCs Grass86 ; later called ‘stochastic complexities’ in Young ). The fact that the algorithm proposed in Zambella is much more general was stressed in arXiv , where also a number of other points were clarified.
More precisely, we consider a finite stationary ergodic irreducible first order Markov process $M$. Here, “finite” means that it has a finite number $N$ of states, in the following called $s_{t}$ with values numbered simply by $i",\;i=1\ldots N$. The restriction to first order is trivial, as any higher order Markov chain can be re-written as a first order process by re-labelling the states. Irreducibility is essential since it implies (by the Frobenius-Perron-theorem) that there is a unique stationary state. The probability to be in state $i$ at (discrete) time $t$ is denoted as $P_{i,t}$. The stationary state is characterized by the probabilities $P_{i,\infty}$. The transition matrix is $T_{ji}={\rm prob}\{s_{t+1}=j|s_{t}=i\}$. The time evolution of the probabilities is
@@
The optimally inferred states of the hidden dynamics based on the observation process were called ‘mixed states’ in Jurgens1 ; Jurgens2 ; Jurgens3 (papers I-III), and will be denoted here as $\hat{\bf P}_{t}(x_{1}x_{2}\ldots x_{t})$. Based on them, the optimal forecast of the observed process is $\hat{p}_{t+1}(x|x_{1}x_{2}\ldots x_{t})=\hat{p}_{t+1}(x_{1}x_{2}\ldots x_{t},x)/\hat{p}_{t}(x_{1}x_{2}\ldots x_{t})$ with
@@
Equations (4) and (6) are essentially the same as Eq.(3.4) in Zambella , and are today well known as ‘forward algorithm’ Zucchini . They hold also for ‘observable operator models’ (OOMs) Jaeger ; Spanczer with discrete time and discrete values (in which the positivity requirements made for HMMs are dropped), and even for ‘matrix product states’ in statistical physics Derrida ; Klumper ).
@@
As shown in Zambella (see also arXiv ), the set of ‘mixed states’ becomes precisely the set of causal states introduced in Grass86 when minimized. Minimization means that two vectors $\hat{\bf P}_{t}(x_{1}\ldots x_{t})$ and $\hat{\bf P}_{t}(x^{\prime}_{1}\ldots x^{\prime}_{t})$ are identified if they lead, via Eq.(4), to the same $\hat{p}_{t+k}(x)\;\forall k\geq 1$. No such minimization was done in I, which is why no FCs were estimated there and no precise estimates of entropy. It was also shown in Zambella that causal states do in general not correspond to elements of a partitioning of the space of complete histories (in contrast to what was claimed e.g. in Crutch_physica ; Shalizi ; Feldman ; Loomis ), at least when transient parts of the $\epsilon$-machine” are not cut off. Cutting them off (as proposed in most recent papers by Crutchfield et al.) is not a viable alternative, as in some cases (called ‘totally recurrent graphs’ in arXiv ) none of the nodes of the $\epsilon$-machine” correspond to an element of a partitioning of (complete, infinite) histories Zambella ; arXiv . An example of the latter is e.g. ‘rule 22’ Zambella ; arXiv , another example is ‘rule 4’ arXiv .
@@
A main point in I - III was that no such minimization can in general be expected (and FCs are infinite), if the HMM depends on irrational real valued parameters, such that all different observed sequences lead to different optimal forecasts. It was already pointed out in Grass-Helv ; Grass-Nat that even rather simple systems can lead to infinite FC. On the other hand, symmetries of the ${\bf T}^{(x)}$ can lead to non-trivial minimizations, even if the ${\bf T}^{(x)}$ depend on irrational parameters (an instance where some optimal forecasts do not depend on the history at all, but FC is infinite nevertheless, will be be discussed below). No entropy estimates with precision close to these in Zambella were made in I, and no estimates of FCs at all. The main reason is that no minimizations of the set of ‘mixed states’ were performed.
It is claimed in I-III that the set of mixed states is in general not countable, if irrational parameters are involved, but this is wrong. It is always countable. Let us discuss this very important point somewhat more in detail, because it is also related to the very definition of $\epsilon$-machines” and of causal states, and to the question whether the latter correspond to partitionings of history space or not — and because it is also responsible for the slow progress made in ”computational mechanics”, whose state of art  [^1] Ref. Zambella still contains the only precise estimates of FC for non-trivial models, Grass-Nat gives the only such estimates for what is called “set complexity” (SC) in Grass86 , and Zambella contains the only discussion of what are called “totally recursive” systems in arXiv . is in many ways still represented by Zambella ; Grass-Nat .
All these constructs were introduced Grass86 ; Young with the goal of forecasting, more precisely with the goals of performing actual forecasts and quantifying the difficulties involved  [^2] According to its title, a main goal in Young was also inferring complexity, but this was never seriously persued, except for Strelioff , where however only extremely simple models were treated. The difficulties in inferring structural complexity are discussed in arXiv .. In such an enterprise, one cannot deal with histories extending into the infinite past, since they would a.s. not be finitely describable (except in trivial models). As in the theory of automata and formal languages Hopcroft and as discussed above, one must thus assume that observations started at some finite time, and one thus deals with finite histories. The use of infinite histories in constructing $\epsilon$-machines”, as done e.g. in Crutch_physica ; Shalizi ; Feldman ; Loomis , is thus based on a fundamental misunderstanding. From this follows directly that complete $\epsilon$-machines” must include start nodes (as, e.g. in Young ), that their nodes (the causal states) are countable, and that causal states correspond to equivalence classes of finite histories but to coverings of the set of infinite histories. The assumption in I-III and in Crutch_physica ; Shalizi ; Feldman ; Loomis that the definition of causal states is based on infinite histories makes the theory more elegant but much less useful – and is manifestly wrong for “totally recurrent” models Zambella ; arXiv .
Using infinite histories would lead to $\epsilon$-machines without transient parts and start nodes. This looks like a simplification, but makes a great technical problem: Using start nodes was essential in calculating entropies and forecasting complexities in Zambella , ‘set complexities’ in Grass-Nat , and “transient informations” Feldman in arXiv .
It is not clear when the transition happened from $\epsilon$-machines with start nodes and transient parts as in Young (and thus with causal states corresponding to finite histories) to infinite histories and $\epsilon$-machines without start nodes as in Crutch_physica ; Shalizi ; Feldman ; Loomis . Anyhow, it was never mentioned.
@@
Thus the set of “mixed” (i.e., of correct causal) states is always countable by construction. What is correct, however, is that the closure of the set of mixed states (which includes states corresponding to infinite memory) is in many cases uncountable, just as the rational numbers are countable, while the reals (their closure) are not (a non-countable set could be obtained by using any arbitrary starting vector ${\bf P}_{0}$ instead of ${\bf P}_{0}={\bf P}_{\infty}$, but the resulting forecasts would then not be optimal). This non-countability of the closure is, however, not restricted to HMMs with real-valued parameters (as claimed in I - III), but holds also for many of the HMMs studied in Zambella . The main difference between simple and complex models is often not the cardinality of the sets of inferences, but is the Hausdorff dimension of the invariant measure (the Blackwell measure Blackwell ) supported by it. It was zero in all models studied in Zambella , although the box-counting dimensions were non-zero in many cases.
@@
The failure of distinguishing sets from their closures is also responsible for most of II being obsolete. As remarked already in Zambella , the set of mixed (= correct causal) states is very similar to an iterated function system (IFS), with a difference largely overlooked in I - III: While the maps ${\bf T}^{(x)}$ are chosen randomly (maybe with non-uniform probabilities) in an IFS, their choice is in the present case constrained by the (in general non-trivial) grammar of the sequence $\{x_{t}\}$. Thus theorems about IFSs cannot be taken over blindly  [^3] Non-trivial grammatical restrictions do not seem to be responsible for the ‘resetting strings’ found in some of the models studied in Zambella , which are finite strings $R=x_{1}\ldots x_{n}$ of observed symbols for which ${\bf T}^{(R)}\equiv{\bf T}^{(x_{n})}\ldots{\bf T}^{(x_{1})}$ maps any $\bf P$} onto $\bf P$$_{\infty}$. They correspond to catastrophic forgettings of the internal state and are then responsible for the finite Hausdorff dimension of the Blackwell measure; in a systematic search, we now found that this phenomenon occurs in 108 of the 256 models studied in Zambella . For the model called ‘rule 22’ in Zambella . the shortest resetting string is e.g. ‘01000010000’ (the string given in Zambella is an extension of this).. Nevertheless, the open set condition (OSC) Moran ; Peres-Rams should hold in many cases. But its application is much less trivial than assumed in II, where the (non-)overlapping of the sets was confused with the (non-)overlapping of their closures. That such a confusion can be fatal is maybe best illustrated by the ‘fat baker’s transformation’ Alexander . Another example where it leads to a wrong conclusion, proposed originally in paper III, is discussed below.
@@
If minimization of the set of mixed states had been used, the study of the Hausdorff dimension of the set of causal states in II (called ‘statistical complexity dimension’ there), would have been very useful. But although this minimization was alluded to repeatedly in I - III, it was never actually done.
@@
Indeed, this can be defined for general stationary ergodic processes, not only for HMM. All what is needed is that one can define a sequence of approximate HMMs or OOMs, to which one can apply Eqs. (4,6). Minimizing $h_{C}-h$ can be used for most accurate model (or parameter) inference Zucchini , but minimizing both $h_{C}-h$ and $C$ at fixed $\alpha$ could be useful for finding models / parameters / approximations that are both accurate and relatively easy to forecast.
@@
Finally, Ref. Jurgens3 (paper III) is devoted to the calculation of the ‘ambiguity rate’ $h_{a}$ of a process, a notion supposedly introduced by Shannon in Shannon as a quantity opposed to ‘equivocation’. In fact, ‘ambiguity’ is mentioned exactly once in Shannon , but as a synonym for equivocation. The concept as defined in III is the rate at which information about the past can be discarded, and yet prediction of the future is optimal. If $h_{a}<h$, this would mean that less information is dismissed than new information is obtained, i.e. that FC (the information stored in an optimal forecasting scheme) diverges with time. In the opposite case, i.e. when $h_{a}=h$, it was claimed in III that “the associated $\epsilon$-machine’s internal causal-state process is stationary.” This is wrong, as demonstrated by the counter-examples given in Zambella . There it was shown that FC is finite in all considered cases (i.e., $h_{a}=h$), but yet the forecasting process is not stationary in cases like ‘rule 22’, because of the slow convergence of FC.
@@
When each observed sequence leads to a different forecast, it is clear that no information can be dismissed in an optimal forecast, and thus $h_{a}=0$. In some cases, the situation is a bit more subtle. Consider, e.g., the example in Sec. 7B of III for control parameter values $x=0.25,\alpha=0.5$, which was treated in detail in appendix B of III. In this model, there are three emitted symbols $\triangle,\Box,\circ$. One finds that always $\hat{p}_{t}(\Box)=1/4$, while $\hat{p}_{t}(\triangle)$ and $\hat{p}_{t}(\circ)$ depend non-trivially on the observations at times $1,\ldots t-1$. Indeed, no two different histories lead to the same $\hat{p}_{t}(\triangle)$ and $\hat{p}_{t}(\circ)$ (at least for $t<12$; for larger $t$ one would need extended precision arithmetic to avoid numerical problems). For each $t$, the largest values of $\hat{p}_{t}(\triangle)$ and $\hat{p}_{t}(\circ)$ coincide, and so do the smallest values. The difference between the two increases with $t$ and tends to $0.07020(1)$ for $t\to\infty$, which shows clearly that the history dependence does not vanish asymptotically. Thus, no information can be dismissed if all forecasts are to be optimal, and $h_{a}=0$ — while $h_{a}=0.4499$ bits is quoted in III. Obviously, this wrong value was obtained because $h_{a}$ was calculated in III using the open set condition, and overlaps between sets were confused with overlaps between their closures. Since this error was made in all calculations of $h_{a}$, all numerical results of III are either trivial or obsolete.
Acknowledgements: I thank Sarah Marzen and Paul Riechers for extremely useful discussions.